OpenStudy (anonymous):

y=x-4 1.(2,6) 2.(6,2) 3.(-2,6) 4.(-6,2)

jimthompson5910 (jim_thompson5910):

What's your question?

OpenStudy (anonymous):

the question is what is y=x-4

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

that's all it asks?

jimthompson5910 (jim_thompson5910):

it seems odd they also list a bunch of points

OpenStudy (anonymous):

yes that's basically it

jimthompson5910 (jim_thompson5910):

are they asking anything like if they are solutions to the equation?

OpenStudy (anonymous):

yes they are asking if any of those ordered pairs are the solution to y=x-4

jimthompson5910 (jim_thompson5910):

Let's test (2,6)

jimthompson5910 (jim_thompson5910):

what is x and y if we have the point (2,6) ?

OpenStudy (anonymous):

x=2 and y=6

jimthompson5910 (jim_thompson5910):

y=x-4 y=2-4 ... replace x with 2 (since x = 2) 6 = 2-4 ... replace y with 6 (since y = 6) 6 = -2 ... this equation is FALSE

OpenStudy (anonymous):

ok so whats next?

jimthompson5910 (jim_thompson5910):

since plugging in (x,y) = (2,6) into y=x-4 leads to a false equation, this means (2,6) is NOT a solution to y=x-4 visually, the point (2,6) is NOT on the line y=x-4

OpenStudy (anonymous):

how about (-2,6)

jimthompson5910 (jim_thompson5910):

plug (x,y)=(-2,6) into y=x-4 and tell me what you get

OpenStudy (anonymous):

I actually did (-6,2) y=x-4 2=-6-4 2=2

jimthompson5910 (jim_thompson5910):

-6-4 is not equal to 2

jimthompson5910 (jim_thompson5910):

-6-4 is equal to -10

jimthompson5910 (jim_thompson5910):

so 2 = -6-4 2 = -10 which is false, so (-6,2) is not a solution to y = x-4